OpenStudy (anonymous):

What is the slope of the line that contains the points (2, -6) and (-3, 1)? a) -7/5 b) -1 c) -5/7 d) 1 please help !

OpenStudy (anonymous):

(y-y1)*(x2-x1)=(x-x1)*(y2-y1) here (x1,y1),(x2,y2) are given points then write them in the form of y=mx+c slope is m

OpenStudy (anonymous):

or formula slope=(y2-y1)/(x2-x1)

jimthompson5910 (jim_thompson5910):

use the slope formula \[\Large m = \frac{y_{2} - y_{1}}{x_{2} - x_{1}}\]

jimthompson5910 (jim_thompson5910):

In this case (x1,y1) = (2, -6) (x2,y2) = (-3, 1)

OpenStudy (anonymous):

alright , thank you ! :)